Jim realizes that he screwed up at the last Engineering
Competition so he decides to build a rocket to impress
Cindy. The rocket is fired into the air and reaches its highest
point 4 seconds later.

A) With what velocity was the rocket fired?

B) How high did the rocket go up?
1
Expert's answer
2020-10-15T03:13:20-0400

a)


"v=gt=(9.8)(4)=39\frac{m}{s}"

b)


"h=0.5gt^2=0.5(9.8)(4)^2=78\ m"